what is the measure of each angle if the measure of one angle is 5 times the measure of it complement
Answer by ad_alta(240)

Let 'a' be the angle and 'c' its complement. Then a=5(c) and since c=90-a we have a=5(90-a). That gives a=75 and c=15.